OpenStudy (anonymous):

WILL GIVE FAN AND MEDAL Find an equation for the nth term of a geometric sequence where the second and fifth terms are -8 and 512, respectively.

OpenStudy (anonymous):

@TrojanPoem Would you happen to know how to solve this?

OpenStudy (trojanpoem):

x_{2} = -8 x_{5} = 512 ar^2-1 = -8 ar^5-1 = 512 ar = -8 ar^4 = 512 divide both eqn r^3 = 64 r = 4 a = -2 eqn = -2 (4)^(n-1)

OpenStudy (anonymous):

so 2(-4)^n-1 makes sense?